OpenStudy (anonymous):

what is the SLOPE OF A LINE 3(x * x) + 10xy + 3 ( y*y ) - 15x - 21y + 18 = 0

OpenStudy (anonymous):

you can find the slope by finding teh derivative. i think you should use implicit differentiaton. it will be: 6x+10y+10(dy/dx) + 6y(dy/dx) -15-21(dy/dx) = 0 (dy/dx) (10+6y-21) = 15-6x-10y dy/dx = (15-6x+10y)/(10+6y-21) = m do you know any points on the line? i dont think i can find the slope if we dont know any point on the line